evaluate ∫ (1 – cos x) / (1 + cos x) dx

evaluate
∫  (1 – cos x) / (1 + cos x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

I = \int \frac{1-cosx}{1+cosx}dx
I = \int \frac{2sin^{2}\frac{x}{2}}{2cos^{2}\frac{x}{2}}dx
I = \int tan^{2}\frac{x}{2}dx
I = \int (sec^{2}\frac{x}{2}-1)dx
I = \frac{tan\frac{x}{2}}{1/2}-x + c
I = 2tan\frac{x}{2}-x + c